OpenStudy (anonymous):

1. General Electric won a bid to supply jet engines with a total list price of $600 million to China Airlines. Assuming a trade discount of 20 percent, what did General Electric receive?

OpenStudy (anonymous):

on $ 100 the discount would be $ 20 therefore on $ 1 the discount would be 20/100 = 1/5 on $ 600 million the discount = 1/5x $ 600 million = $ 120 million Money Received by General Electric = $(600-120) million = $480 million

OpenStudy (anonymous):

Oh I get it now thank you very much. ;-)